# RAISE: Enhancing Scientific Reasoning in LLMs via Step-by-Step Retrieval

### v2

@@ Abstract
Scientific reasoning requires not only long-chain reasoning processes, but also knowledge of domain-specific terminologies and adaptation to updated findings. To address these challenges in scientific reasoning, we introduce RAISE, a step-by-step retrieval-augmented framework which retrieves logically relevant documents from an in-the-wild corpus. RAISE is divided into three steps: problem decomposition, logical query generation, and logical retrieval. We observe that RAISE consistently outperforms other baselines on scientific reasoning benchmarks. We analyze that unlike other baselines, RAISE retrieves documents that are not only similar in terms of the domain knowledge, but also documents logically more relevant.
RAISE: Enhancing Scientific Reasoning in LLMs via
Step-by-Step Retrieval
Minhae Oh1, Jeonghye Kim2, Nakyung Lee1, Donggeon Seo3, Taeuk Kim3, Jungwoo Lee1, 1Seoul National University, 2KAIST, 3Hanyang University
@@ 1 Introduction
Recently, large language models (LLMs) have demonstrated remarkable capabilities in scientific domains that require intricate reasoning and the integration of diverse knowledge sources Zhang et al. (2024c, a); Prabhakar et al. (2025); Rueda et al. (2025). In contrast to many other domains, scientific reasoning is especially demanding due to the need for advanced reasoning skills, knowledge of specialized terminology, and awareness of continually evolving knowledge.
To enhance the capability of LLMs to solve scientific reasoning problems, two approaches are widely used: (1) step-wise reasoning, and (2) retrieval-augmented generation. For step-wise reasoning, previous studies Wei et al. (2022); Zhao et al. (2024a); Jin et al. (2024); Lai et al. (2024); Yao et al. (2023) have shown that solving complex problems through a step-by-step process is more effective than attempting to solve them in a single step, mirroring how humans approach challenging problems. For retrieval-augmented generation (RAG), it is known that retrieving relevant external knowledge can mitigate hallucinations and help models adapt to newly emerging information, which is especially important in the scientific domain Lewis et al. (2020b); Bang et al. (2025); Zhong et al. (2025); Xiong et al. (2024).
@@
To utilize the benefits of both worlds, recent works have adopted RAG for step-wise reasoning across various applications Jeong et al. (2024); Zhang et al. (2025a). However, most studies focus on relatively simple multi-hop question answering tasks or non-scientific domains Guan et al. (2025); Jin et al. (2025); Wang et al. (2025), or assume the availability of task-specific retrieval corpora, which are more difficult to obtain than widely accessible web-based sources such as Wikipedia Zhong et al. (2025); Qian et al. (2023); Anand et al. (2023). Solving challenging scientific reasoning tasks,such as graduate-level biology or chemistry, using an in-the-wild corpus is difficult since merely retrieving superficial knowledge is insufficient. Instead, the retrieved information should contain relevant logical connections needed to solve the problem Rueda et al. (2025). Moreover, the knowledge required for each intermediate step can vary significantly even within the same problem. Without considering the evolving information needed for each reasoning process, RAG might even deteriorate the downstream task performance. The question of what to search for and how to retrieve the appropriate external knowledge for each step when solving scientific reasoning tasks is underexplored.
Figure 1: Overview of RAISE. RAISE is divided into three steps: (1) Problem Decomposition, (2) Logical Query Generation, and (3) Logical Retrieval. Given an scientific reasoning question, we decompose the problem into multiple subquestions and search queries. Then, we generate a logical query that contains logic-enriched contexts. Finally, we retrieve relevant documents from in-the-wild corpus for each subquestion and condition the documents to assist the step-by-step reasoning for scientific reasoning problems.
@@
To address these challenges, we introduce RAISE (Step-by-Step Retrieval-Augmented Inference for Scientific rEasoning), a retrieval-augmented framework tailored for step-wise scientific reasoning. Our framework is divided into three steps: (1) problem decomposition, (2) logical query generation, and (3) logical retrieval. For problem decomposition, LLMs break down the original question into subquestions along with search queries that can assist in solving each reasoning step. Then, for each search query, LLMs transform it into a logical query—one that includes the underlying logical explanations. Since logical queries contain more information about what external knowledge is needed to solve the problem, they enable retrieval based on logical rather than superficial domain similarity. Finally, we perform retrieval separately for each step, obtaining documents that are logically relevant to the particular step. Instead of assuming task-relevant or well-curated retrieval source, such as question-answer pool of relevant domains, we retrieve from in-the-wild source such as Wikipedia, which enables applying to challenging real-world scenarios.
We evaluate our framework on three challenging scientific reasoning benchmarks: GPQA Rein et al. (2024), SuperGPQA Team et al. (2025), and MMLU Hendrycks et al. (2021); Wang et al. (2024b). Across these datasets, RAISE results in consistent performance improvement compared to various baseline approaches that include applying RAG or problem decomposition. This indicates that our framework effectively retrieves relevant documents that are needed to solve scientific reasoning tasks step-by-step. Through further analysis, we observe that RAISE retrieves more logically relevant documents for each step compared to baselines, highlighting that logical similarity is crucial for these challenging reasoning tasks in addition to domain similarity.
@@ 2 Preliminary
[…]
@@
Figure 2: Examples comparing query generation methods (Step-Back, HyDE, and RAISE) for the same subquestion.
@@ 3 RAISE
We propose RAISE (Step-by-Step Retrieval-Augmented Inference for Scientific rEasoning), a retrieval-augmented generation framework for scientific reasoning designed to support multi-step reasoning through fine-grained, step-aware retrieval. Rather than retrieving documents based on the entire problem as a single query, RAISE decomposes the problem into intermediate reasoning steps and performs targeted retrieval for each. The overview of RAISE is provided in Figure 1 and Algorithm 1 in Appendix A. The method consists of three main stages:
  1. (1)
    Problem Decomposition, where the original question $x$ is broken down into subquestions $r_{1},\dots,r_{n}$ along with corresponding search queries $q_{1},\dots,q_{n}$ that are intended to retrieve relevant knowledge for each step;
  2. (2)
    Logical Query Generation, where each search query $q_{i}$ is further enriched into a logic aware explanation $q_{i}^{*}$, which serves as the final retrieval query; and
  3. (3)
    Logical Retrieval, where external knowledge $D_{i}$ is retrieved for each subquestion $r_{i}$ and used to generate the subanswer $a_{i}$.
@@
RAISE begins with the decomposition of a scientific reasoning problem $x$ into a sequence of intermediate subquestions $\{r_{i}\}_{i=1}^{n}$, each corresponding to a distinct reasoning step. This planning process is performed by prompting a language model $P_{\theta}$ with $x$ and prompt $p_{1}$ to identify the logical structure required to solve the problem. Formally, we generate:
$\{(r_{i},q_{i})\}_{i=1}^{n}\sim P_{\theta}(\cdot\mid x,p_{1}).$ (1)
While recent work has increasingly explored problem decomposition, many retrieval-based approaches still operate at the problem level, retrieving information using a single query without explicitly modeling intermediate reasoning steps. In contrast, this stage produces a structured sequence of subquestions that reflects the multi-step nature of scientific reasoning.
Along with each subquestion $r_{i}$, a corresponding search query $q_{i}$ is generated to guide retrieval. These queries are not used directly for retrieval but rather serve as an initial query for the next stage. As a result, this stage outputs subquestion-query pairs $\{(r_{i},q_{i})\}_{i=1}^{n}$, forming the basis for step-wise retrieval and generation.
@@
In the second stage, each initial search query $q_{i}$ and its corresponding subquestion $r_{i}$ are jointly used to generate a logically enriched logical query $q_{i}^{*}$. Search queries $q_{i}$ are short key phrases, where the logical process is absent. Therefore, using $q_{i}$ for retrieval will focus on domain similarity instead of logical similarity. Instead, this step constructs a logically coherent query that aligns with the reasoning intent of the subquestion. The model is prompted with both $q_{i}$ and $r_{i}$, along with a reformulation prompt $p_{2}$ (e.g., add essential scientific or mathematical explanation the search result would contain):
$q_{i}^{*}\sim P_{\theta}(\cdot\mid q_{i},r_{i},p_{2})$ (2)
Using only the subquestion $r_{i}$ to generate a query often results in noisy or overly specific formulations that include problem-specific variables or conditions irrelevant to general knowledge retrieval. On the other hand, using only the initial query $q_{i}$ lacks sufficient context to capture the intended reasoning path. By combining both, we enable the model to synthesize a logical query that is purpose-driven and aligned with the knowledge required to answer $r_{i}$. Even if the reformulated query $q_{i}^{*}$ contains factual inaccuracies, it tends to retrieve passages from a corpus $\mathcal{C}$ that are logically relevant and supportive of the reasoning required for solving the original problem. Figure 2 presents example queries generated by RAISE, Step-Back+RAG, and HyDE, illustrating RAISE’s ability to generate logical queries that are well-aligned with the reasoning intent.
@@
The reformulated query $q_{i}^{*}$ is used to retrieve external knowledge:
$D_{i}=\mathcal{R}(q_{i}^{*},\mathcal{C},k)$ (3)
[…]
After retrieval, for each subquestion $r_{i}$, the model predicts its solution $a_{i}$. Importantly, the prediction is conditioned not only on the retrieved documents $D_{i}$, but also on the original question $x$ and the previously answered subquestions $\{(r_{j},a_{j})\}_{j=1}^{i-1}$. We write:
$a_{i}\sim P_{\theta}(\cdot\mid x,\{(r_{j},a_{j})\}_{j=1}^{i-1},r_{i},D_{i},p_{3})$ (4)
Finally, all subanswers are aggregated to predict the final answer $y$:
$y\sim P_{\theta}(\cdot\mid x,\{(r_{i},a_{i})\}_{i=1}^{n},p_{4})$ (5)
This step-wise approach ensures that the model constructs its final answer based on grounded and explicitly structured intermediate reasoning.
@@
Overall
science-
hard
science-
middle
engineering-
hard
Chemistry
Biology
College Chemistry
@@
RAISE
51.01
(+9.8%)
10.05
(+33.3%)
19.60
(+25.8%)
10.55
(+5.0%)
28.36
(+11.5%)
59.27
(+2.2%)
51.00
(+3.0%)
Table 1: Comparison of various reasoning strategies across GPQA, SuperGPQA, and MMLU. The underscore indicates the highest score among baselines, while the boldface indicates the highest score overall. The value in parentheses below the RAISE score represents the performance gain compared to the best baseline. RAISE consistently outperforms other approaches for scientific reasoning benchmarks.
4 Experiments
@@ 4.1 Experimental Setup
4.1.1 Datasets.
To evaluate our method in settings that require multi-step scientific reasoning, we select three datasets focused on high-level STEM and professional reasoning. We provide more details on the datasets in Appendix B.
  • GPQA (Graduate-level Physics Question Answering) Rein et al. (2024): A collection of graduate-level physics, biology, and chemistry questions requiring multi-step reasoning and access to external scientific fact. We use this dataset as our primary benchmark.
  • SuperGPQA Team et al. (2025): This dataset evaluates the knowledge and reasoning abilities of LLMs across 285 graduate-level disciplines. We focus on three subsets that are challenging enough and require step-by-step reasoning in the science domain: science-hard, science-middle, and engineering-hard.
  • MMLU Hendrycks et al. (2021); Wang et al. (2024b): We focus on a subset of science-related tasks in MMLU-Pro, specifically chemistry and biology, which are well-suited for evaluating step-by-step reasoning with specialized knowledge. Given the lower overall difficulty of MMLU-STEM compared to MMLU-Pro, we additionally include its chemistry category as a lightweight validation setting.
4.1.2 Retriever Configuration.
We adopt Dense Passage Retrieval (DPR) Karpukhin et al. (2020) as our retriever, using the publicly available model trained on the Natural Questions (NQ) dataset Kwiatkowski et al. (2019).
4.1.3 Language Model.
For GPQA, which is the primary focus of our study, we use Mistral Small 3.1-Instruct-2503 Mistral AI (2025), a 24B state-of-the-art small-scale open-sourced LLM that achieves strong performance on scientific reasoning tasks. For the remaining tasks, we use the smaller LLaMA 3.1-8B model Grattafiori et al. (2024) due to computational constraints, applying it to SuperGPQA, and MMLU. We provide detailed hyperparameters and experiment setups in Appendix C.
4.1.4 Baselines.
To assess the importance of multi-step reasoning and step-aware retrieval, we conduct experiments with four groups of baselines: Direct Reasoning, Direct Reasoning with RAG, Decomposed Reasoning, and Decomposed Reasoning with RAG.
Direct Reasoning solves problems in a single step without decomposing them into subquestions, and includes Chain-of-Thought (CoT) Wei et al. (2022), which generates intermediate reasoning steps through prompting. A retrieval-augmented variant, Direct Reasoning with RAG, includes CoT+RAG Lewis et al. (2020b), which retrieves documents using the original question as the query.
Decomposed Reasoning methods break down the original problem into subquestions and solve them sequentially. This includes Least-to-Most Zhou et al. (2023), which solves each subquestion directly, and Step-Back Zheng et al. (2024), which applies principle extraction at the subquestion level. A retrieval-augmented variant, Decomposed Reasoning with RAG, retrieves evidence for each subquestion and solves them step-by-step. This group includes Least-to-Most+RAG Lewis et al. (2020b), Step-Back+RAG Zheng et al. (2024), and HyDE Gao et al. (2023a), with the latter two enhancing retrieval relevance through query reformulation, making them strong baselines for comparison.
@@ 4.2 Main Results
As shown in Table 1, our proposed method, RAISE, consistently outperforms all baseline reasoning strategies across three benchmark datasets of varying difficulty: GPQA, SuperGPQA, and MMLU, achieving an average performance improvement of 13% over the best baseline scores. Unlike other baselines whose performance varies depending on the dataset’s difficulty or type, RAISE consistently demonstrates robust performance and outperforms them across different domains, types, and levels of difficulty. Specifically, on challenging GPQA, RAISE achieves a score of 51.01, outperforming the next-best method (HyDE) by 9.8%.
@@
To assess the effectiveness of our logical query generation, we compare RAISE with three RAG-based decomposed reasoning baselines that differ in how they construct retrieval queries. Least-to-Most+RAG uses the subquestion itself as the query, Step-Back+RAG abstracts a general principle from the subquestion, and HyDE generates a hypothetical answer to use as the retrieval query. RAISE consistently outperforms all baselines across benchmarks, demonstrating the advantage of generating logically grounded queries that better align with the reasoning required to solve each subquestion. These results confirm that RAISE’s queries go beyond retrieving documents that are merely domain-relevant or superficially similar, enabling access to knowledge that is logically aligned with the problem-solving process.
Unlike RAISE, decomposed reasoning methods do not always yield better performance, particularly for smaller LLMs that lack logically relevant knowledge Hosseini et al. (2024); Xu et al. (2025). While prior work Zhou et al. (2023) using decomposition focused on proprietary large models, our study targets open-source smaller LLMs. Without sufficient background knowledge, even well-structured subquestions can lead to hallucinations or unsupported reasoning, making decomposition potentially counterproductive unless relevant documents are retrieved.
We also observe cases where models without retrieval outperform their retrieval-augmented variants. For example, on the SuperGPQA engineering-hard subset, Least-to-Most+RAG (8.04) performs worse than Least-to-Most (10.05). Although retrieval is intended to improve performance by providing additional information, naive retrieval can introduce noise, especially in open-domain settings where irrelevant content is common. Conventional RAG methods often retrieve documents that are only superficially related or worse, focus on irrelevant variables or distractors rather than the core principles required for reasoning. These results highlight the importance of retrieving logically relevant knowledge and further support the effectiveness of our approach.
4.2.2 Applying RAISE to various LLMs.

### v3

@@ Abstract
Scientific reasoning requires not only long-chain reasoning processes, but also knowledge of domain-specific terminologies and adaptation to updated findings. To deal with these challenges for scientific reasoning, we introduce RAISE, a step-by-step retrieval-augmented framework which retrieves logically relevant documents from in-the-wild corpus. RAISE is divided into three steps: problem decomposition, logical query generation, and logical retrieval. We observe that RAISE consistently outperforms other baselines on scientific reasoning benchmarks. We analyze that unlike other baselines, RAISE retrieves documents that are not only similar in terms of the domain knowledge, but also documents logically more relevant. Codes are available at https://github.com/minhaeoh/RAISE.
@@ 1 Introduction
Large language models (LLMs) have shown strong potential for scientific reasoning, which demands advanced reasoning skills, domain-specific terminology, and up-to-date knowledge [45, 42, 27, 30]. Two common strategies are step-wise reasoning, which solves complex problems through structured intermediate steps [37, 47, 12, 18, 41], and retrieval-augmented generation (RAG), which mitigates hallucinations by providing external evidence [21, 1, 50, 39]. Recent work combines them, but often targets simpler multi-hop QA or assumes curated, task-specific corpora [9, 44, 5, 11, 34], unlike open-domain sources such as Wikipedia. Solving challenging scientific reasoning tasks, such as graduate-level biology or chemistry, using an in-the-wild corpus is difficult since merely retrieving superficial knowledge is insufficient. Instead, the retrieved information should contain relevant logical connections needed to solve the problem [30]. Moreover, the knowledge required for each intermediate step can vary significantly even within the same problem. Without considering the evolving information needed for each reasoning process, RAG might even deteriorate the downstream task performance. The question of what to search for and how to retrieve the appropriate external knowledge for each step when solving scientific reasoning tasks is underexplored.
Figure 1: Overview of RAISE. RAISE is divided into three steps: (1) Problem Decomposition, (2) Logical Query Generation, and (3) Logical Retrieval.
@@
To address these challenges, we introduce RAISE (Step-by-Step Retrieval-Augmented Inference for Scientific rEasoning), a retrieval-augmented framework tailored for step-wise scientific reasoning. Our framework consists of three stages: (1) problem decomposition, where LLMs break down the original question into subquestions along with search queries; (2) logical query generation, which reformulates each search query into a logic-enriched query that captures the reasoning needed to solve the subquestion; and (3) logical retrieval, which retrieves step-specific documents from an open-domain corpus, ensuring the retrieved information is logically relevant rather than superficially domain similar. Instead of assuming task-relevant or well-curated retrieval source, such as question-answer pool of relevant domains, we retrieve from in-the-wild source such as Wikipedia, which enables applying to challenging real-world scenarios. Evaluated on GPQA, SuperGPQA, and MMLU, RAISE consistently outperforms baselines using either RAG or problem decomposition alone, demonstrating its ability to retrieve step-specific, logically relevant information essential for solving complex scientific reasoning tasks. While this work focuses on scientific reasoning benchmarks, similar challenges arise in mathematical reasoning, which also requires precise, multi-step logical inference.
@@ 2 Preliminary
Top-$k$ documents are retrieved, and a generator (e.g., BART [20], T5 [28]) produces $y$ based on both $x$ and $d_{i}$.
@@ 3 RAISE
We propose RAISE (Step-by-Step Retrieval-Augmented Inference for Scientific rEasoning), a retrieval-augmented generation framework for scientific reasoning designed to support multi-step reasoning through fine-grained, step-aware retrieval. The method consists of three main stages: (1) Problem Decomposition, (2) Logical Query Generation, (3) Logical Retreival. The overview of RAISE is provided in Figure 1 and Algorithm 1 in Appendix A.
@@
RAISE decomposes the problem $x$ into subquestions $r_{1},\dots,r_{n}$ with corresponding search queries $q_{1},\dots,q_{n}$, forming a structured sequence for step-wise retrieval, unlike conventional single-query approaches. These queries are not used directly for retrieval but rather serve as an initial query for the next stage. As a result, this stage outputs subquestion-query pairs $\{(r_{i},q_{i})\}_{i=1}^{n}$, forming the basis for step-wise retrieval and generation.
@@
In the second stage, each initial search query $q_{i}$ and its corresponding subquestion $r_{i}$ are jointly used to generate a logically enriched logical query $q_{i}^{*}$. Since initial queries $q_{i}$ lacks reasoning context and subquestions $r_{i}$ alone can be noisy or overly specific, neither $q_{i}$ nor $r_{i}$ alone is sufficient for effective retrieval. By combining both, we generate logical queries that better capture the reasoning intent and retrieve logically relevant knowledge for solving each step. The model is prompted with both $q_{i}$ and $r_{i}$, along with a reformulation prompt $p_{2}$. Even if the reformulated query $q_{i}^{*}$ contains factual inaccuracies, it tends to retrieve passages from a corpus $\mathcal{C}$ that are logically relevant and supportive of the reasoning required for solving the original problem. Figure 4 in Appendix C.2 presents example queries generated by RAISE, Step-Back+RAG, and HyDE, illustrating RAISE’s ability to generate logical queries that are well-aligned with the reasoning intent.
@@
External knowledge $D_{i}$ is retrieved for each subquestion $r_{i}$ from in-the-wild corpus $\mathcal{C}$ (e.g., Wikipedia) and used to generate the subanswer $a_{i}$. We retrieve background knowledge for each subquestion using a similarity threshold $T$ to filter irrelevant documents. After retrieval, for each subquestion $r_{i}$, the model predicts its solution $a_{i}$ using $D_{i}$, the original question $x$, and the previous steps. Finally, all subanswers are combined to generate the final answer $y$.
@@
Overall science- hard science- middle engineering- hard (Pro) Chemistry (Pro) Biology (STEM) College Chemistry
@@
RAISE 51.01 (+9.8%) 10.05 (+33.3%) 19.60 (+25.8%) 10.55 (+5.0%) 28.36 (+11.5%) 59.27 (+2.2%) 51.00 (+3.0%)
Table 1: Comparison of various reasoning strategies across GPQA, SuperGPQA, and MMLU. The underscore marks the best baseline, boldface the best overall, and parentheses show RAISE’s gain over the top baseline. RAISE consistently outperforms other approaches for scientific reasoning benchmarks.
4 Experiment
@@ 4.1 Experimental Setup
Datasets.
We evaluate on three scientific benchmarks: GPQA, SuperGPQA, MMLU, which cover graduate-level STEM and professional science tasks that require multi-step scientific reasoning.
Retriever and Language Models.
We adopt Dense Passage Retrieval (DPR) [14] trained on the Natural Questions (NQ) dataset [17] as our retriever. For GPQA, our primary benchmark, we use Mistral Small 3.1-Instruct-2503 [25] (24B), while for SuperGPQA and MMLU we use the lighter LLaMA 3.1-8B model [4] due to computational limits.
Baselines.
To assess the importance of multi-step reasoning and step-aware retrieval, we conduct experiments with four groups of baselines: Direct Reasoning(CoT [37]), Direct Reasoning with RAG(CoT+RAG [21]), Decomposed Reasoning(Least-to-Most [51], Step-Back [49]), and Decomposed Reasoning with RAG. The last group retrieves evidence for each subquestion and solves them step-by-step. This group includes Least-to-Most+RAG [21], Step-Back+RAG, and HyDE [3], with the latter two improving retrieval relevance through query reformulation, making them strong baselines. Further details about datasets, retriever, model settings, and baselines are provided in Appendix B.
@@ 4.2 Main Results
As shown in Table 1, our proposed method, RAISE, consistently outperforms all baseline reasoning strategies across three benchmark datasets of varying difficulty: GPQA, SuperGPQA, and MMLU, achieving an average performance improvement of 13% over the best baseline scores. Unlike other baselines whose performance varies depending on the dataset’s difficulty or type, RAISE consistently demonstrates robust performance and outperforms them across different domains, types, and levels of difficulty. Furthermore, we confirm that these improvements hold across models of different scales, including smaller LLaMA-8B and GPT-4o mini, as shown in Appendix C.1, demonstrating that RAISE’s effectiveness is not tied to a specific LLM architecture or size.
@@
To assess the effectiveness of our logical query generation, we compare RAISE with three RAG-based decomposed reasoning baselines that differ in how they construct retrieval queries. Least-to-Most+RAG uses the subquestion itself as the query, Step-Back+RAG abstracts a general principle from the subquestion, and HyDE generates a hypothetical answer to use as the retrieval query. RAISE consistently outperforms all baselines across benchmarks, demonstrating the advantage of generating logically grounded queries that better align with the reasoning required to solve each subquestion. These results confirm that RAISE’s queries go beyond retrieving documents that are merely domain-relevant or superficially similar, enabling access to knowledge that is logically aligned with the problem-solving process. Qualitative examples further support this finding, as shown in Appendix C.2, where RAISE retrieves passages containing essential scientific mechanisms while conventional RAG often returns vague or unrelated content.
Unlike RAISE, decomposed reasoning methods do not always yield better performance, particularly for smaller open-source LLMs that lack sufficient background knowledge [7, 40]. While decomposition can help structure reasoning, without access to relevant external knowledge, smaller models may produce hallucinations or unsupported intermediate steps, sometimes leading to worse performance than direct reasoning. Moreover, even when retrieval is added, naive RAG can introduce additional noise. In such cases, the retriever may surface superficially related or distracting content rather than the core principles needed for reasoning, which can ultimately harm performance. This highlights the importance of retrieving logically relevant knowledge rather than merely domain-related content, a challenge that RAISE directly addresses.
5 Analysis of RAISE
We also analyze the quality of the retrieved documents. Using both an LLM-as-a-judge and a small-scale human evaluation, we find that RAISE consistently retrieves fewer irrelevant or superficial documents and more passages that directly support reasoning (Appendix C.3). These results confirm that RAISE’s gains stem from retrieving logically aligned knowledge rather than merely domain-related content.
6 Conclusion
We introduce RAISE, a step-by-step retrieval framework for scientific reasoning. We first decompose the problem into multiple subquestions and search queries, and then generate logical queries and retrieve logically relevant documents from in-the-wild corpus. We demonstrate the effectiveness of RAISE on three scientific reasoning benchmarks by comparing with various baselines. Our analysis shows that RAISE retrieves documents that are not only relevant in terms of the domain (e.g. definition of specificalized terminology) but also logically relevant documents for each subquestion, assisting the step-by-step reasoning process required for scientific reasoning. Although our experiments focus on scientific reasoning, the stepwise logical retrieval in RAISE is broadly applicable to other domains such as mathematical problem solving, which also demands precise multi-step inference.
Acknowledgments
This work is in part supported by the National Research Foundation of Korea (NRF, RS-2024-00451435(20%), RS-2024-00413957(20%)), Institute of Information & communications Technology Planning & Evaluation (IITP, RS-2021-II212068(10%), RS-2025-02305453(15%), RS-2025-02273157(15%), RS-2025-25442149(10%) RS-2021-II211343(10%)) grant funded by the Ministry of Science and ICT (MSIT), Institute of New Media and Communications(INMAC), and the BK21 FOUR program of the Education, Artificial Intelligence Graduate School Program (Seoul National University), and Research Program for Future ICT Pioneers, Seoul National University in 2025.
Appendix A RAISE Algorithm
Output: Final answer $y$
Step 1: Problem Decomposition
Generate subquestions and initial queries:
$\{(r_{i},q_{i})\}_{i=1}^{n}\sim P_{\theta}(\cdot\mid x,p_{1})$
for $i=1$ to $n$ do
 Step 2: Logical Query Generation
   Reformulate initial query:
    $q_{i}^{*}\sim P_{\theta}(\cdot\mid r_{i},q_{i},p_{2})$
 Step 3: Knowledge Retrieval
   Retrieve top-$k$ documents:
    $D_{i}=\mathcal{R}(q_{i}^{*},\mathcal{C},k)$
 Step 4: Subquestion Answering
 if $i=1$ then
       $a_{i}\sim P_{\theta}(\cdot\mid x,r_{1},D_{1},p_{3})$
 else
       $a_{i}\sim P_{\theta}(\cdot\mid x,\{(r_{j},a_{j})\}_{j=1}^{i-1},r_{i},D_{i},p_{3})$
   end if
end for
Step 5: Final Answer Composition
Generate final answer using all subanswers:
$y\sim P_{\theta}(\cdot\mid x,\{(r_{i},a_{i})\}_{i=1}^{n},p_{4})$
Algorithm 1 RAISE Inference Procedure
Appendix B Experiment Details
B.1 Dataset Details
GPQA
[29] This dataset consists of physics, biology, and chemistry questions written by domain experts. We use GPQA diamond subset, which consist of 198 high-quality questions selected based on human performance. Specifically, this subset includes questions that both experts answer correctly while the majority of non-experts fail to solve. Each question typically demands multi-step reasoning, precise formula manipulation, and access to external scientific facts (e.g., physical constants, definitions). Due to its alignment with our target setting, GPQA serves as the primary evaluation benchmark throughout our experiments.
SuperGPQA
[32] SuperGPQA is a large-scale benchmark designed to evaluate graduate-level reasoning across 13 disciplines, 72 fields, and 285 graduate-level disciplines. In alignment with the scientific reasoning focus of our work, we select science and engineering domains for evaluation. Each domain is further divided by three difficulty levels(easy, medium, and hard). To reduce computational overhead while maintaining consistency, we randomly sample 199 questions per subset using a fixed seed (42). Specifically, our experiments include 199 examples each from science-hard, science-middle, and engineering-hard subsets.
MMLU
[6, 36] The MMLU benchmark covers a wide range of subjects across multiple domains. For out experiments, we focus on STEM and Professional categories. The STEM contains university-level science and engineering subjects such as college mathematics and computer science, while the Professional category covers specialized fields that typically require professional training or advanced education, including law, medicine, and chemistry. We specifically select three subsets: college chemistry from MMLU-STEM and professional chemistry and biology from MMLU-Pro. These subsets are chosen to evaluate our method’s ability to perform scientific reasoning in both academic and professional contexts involving complex domain knowledge.
B.2 Baseline Details
CoT [37, 16] We apply Chain-of-Thought prompting for direct reasoning, where the model is encouraged to explicitly generate intermediate reasoning steps through prompting (Think step by step).
CoT+RAG
[21] We implement CoT+RAG by combining Chain-of-Thought prompting with retrieval, where the model is prompted to solve the problem step-by-step while also leveraging external knowledge. Specifically, we provide the model with a CoT-style prompt encouraging step-by-step reasoning, alongside the original question and documents retrieved using the original question as the search query.
Least-to-Most
[51] Least-to-Most is a decomposed reasoning strategy that breaks down a complex problem into a sequence of simpler subquestions, which are then solved sequentially without retrieval augmentation. This subquestion decomposition pipeline serves as the foundational structure for other decomposed reasoning methods as well.
Step-Back
(Decomposed reasoning) [49] We implement Step-Back for decomposed reasoning by applying the Step-Back prompting method to each subquestion in a decomposed reasoning framework. While the original Step-Back paper does not cover the application of this method to decomposed subquestions, we extend it for a fair comparison with our approach. Specifically, after decomposing the original question into subquestions, we use the Step-Back prompting strategy to extract a high-level principle for each subquestion, and then provide the subquestion along with its corresponding principle to guide the model’s reasoning.
Least-to-Most+RAG
(Decomposed reasoning with RAG) [23] We implement RAG by first decomposing the original problem into subquestions and then retrieving documents using each subquestion as a query. The retrieved documents are provided to the model along with the corresponding subquestion to support its reasoning.
Step-Back+RAG
(Decomposed reasoning with RAG) [49] We extend the Step-Back prompting strategy to a retrieval-augmented setting for fair comparison with our method. After decomposing the original question into subquestions, we generate a principle abstraction for each subquestion using Step-Back prompting, and use it as a query to retrieve evidence. The retrieved documents are then provided alongside the original subquestion to guide the model’s reasoning.
HyDE
(Decomposed reasoning with RAG) [3] We apply the HyDE approach to each subquestion in a decomposed reasoning framework. For each subquestion, the model first generates a hypothetical answer, which is then used as a query to retrieve supporting documents. The retrieved evidence, together with the subquestion, is provided to the model to support step-by-step reasoning.
B.3 Retriever Configuration
We use the pre-trained DPR encoder from the ’facebook/dpr-question_encoder-single-nq-base’ model [14], which is a BERT-based encoder trained for open-domain question answering. This encoder is trained on the Natural Question (NQ) dataset [17] and is designed to map questions into 768-dimensional dense vector representations for retrieval.
For the retrieval corpus, we use the preprocessed Wikipedia passages provided by ’facebook/wiki_dpr’ [14], a corpus widely used to evaluate DPR-based retrieval models. This corpus is constructed from the December 20, 2018 Wikipedia dump, where each article is split into multiple, disjoint text blocks of 100 words, resulting in approximately 21 million passages. Each passage is accompanied by the title of the wikipedia page it comes from along with DPR embedding.
To enable efficient retrieval over the passage embeddings, we use an exact FAISS index. FAISS (Facebook AI Similarity Search) [13, 2] is widely used library for fast similarity search over dense vectors.
Appendix C Additional Results
C.1 Applying RAISE to various LLMs.